If one side of a parallelogram is 4 cm, other 6 cm, and the diagonal 4√2 cm, 

what is the length of the other diagonal?


1
Expert's answer
2021-06-21T07:32:47-0400

p2 + q2 = 2(a2 + b2 )

where:

p = 4"\sqrt{2}" cm

a = 6 cm

b = 4 cm


Answer = 3 "\sqrt{8}" cm